40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 29 14:39:37 2021, Anchal, Summary, BHD, 1Y1 cleared. IY3 ready for C1SUS2 I/O and FE. RemovedStuff.jpeg1Y1.jpeg1Y3.jpeg
    Reply  Tue Nov 2 16:42:00 2021, Paco, Summary, BHD, 1Y1 rack work 
       Reply  Tue Nov 2 18:54:41 2021, Anchal, Paco, Update, BHD, Optical fibres laid for BHD upgrade 
       Reply  Thu Nov 4 15:03:43 2021, Koji, Summary, BHD, 1Y1 rack work 
          Reply  Mon Nov 8 10:13:52 2021, Paco, Summary, BHD, 1Y1 rack work; Sorensens removed 
             Reply  Mon Nov 8 13:13:00 2021, Koji, Summary, BHD, 1Y1 rack work; Sorensens removed 40m_BHD.png
                Reply  Mon Nov 8 15:29:05 2021, Paco, Summary, BHD, 1Y1 rack work; New power for cameras 
                   Reply  Mon Nov 8 18:42:38 2021, Koji, Summary, BHD, Rack Layout / Power Strips rack_plan.pdf
                      Reply  Tue Nov 9 19:02:47 2021, Anchal, Summary, BHD, 1Y0 Populated and 1Y1,1Y0 powered 
Message ID: 16453     Entry time: Mon Nov 8 10:13:52 2021     In reply to: 16448     Reply to this: 16454
Author: Paco 
Type: Summary 
Category: BHD 
Subject: 1Y1 rack work; Sorensens removed 

[Paco, Chub]

Removed all sorensen power supplies from this rack except for 12 VDC one; that one got pushed to the top of the rack and is still powering the cameras.

ELOG V3.1.3-